INTERNATIONAL CONFERENCE ON DISCRETE DYNAMICS, FINITE AUTOMATA AND COMBINATORICS

23-27 August 1999

URUMQI

Scientific program (Discrete Dynamical Systems, Finite Automata and Combinatorics, Urumqi(Urumchi,Wulumuqi) 1999) :

1) Discrete dynamical systems : symbolic dynamical systems, notions of entropy and complexity, infinite words on a finite alphabet (spectral and ergodic properties), prediction algorithms of a random process, pseudorandom sequences.

2) Finite Automata : Finite automata, substitutions and transductors, autosimilarity, fractal sets, diophantine approximations, systems of numeration, quasicristalline structures and tiling theory, Ising models and automata.

3) Combinatorics : Combinatorial properties of infinite words on a finite alphabet, arithmetical and geometrical properties of systems with low complexity, and other topics on graphs and combinatorics.

Scientific committee Yong-Chuan CHEN (Los Alamos and Nankai Univ.) Teturo KAMAE (Osaka City Univ., Osaka) Yuji ITO (Keio Univ., Yokohama) Xiao-Feng GUO (Xingjiang Univ., Urumqi) Christian Mauduit (Aix-Marseille II, Luminy) Jacques PEYRIERE (Paris XI, Orsay) Zhi-Ying WEN (Tsinghua Univ. and Wuhan Univ., Beijing)

Secretary of the conference Ya-Ru Dang

Venue and dates

The conference will be held in the Xinjiang University (Urumqi, Xinjiang), 23-27, August, 1999.

Call for papers

To submit a paper, you should attach a one-page abstract including the title, author(s), and an abstract of less than 100 words. The deadline for receipt of paper is 30 April, 1999. The abstract could be sent by e-mail in the LaTex file to Prof. GUO Xiao-Feng or WEN Zhi-Ying(see below the e-mail address).

Accommodation 1). Most participants will be housed in the Hotel of Xinjiang University.

If there are not enough rooms in the Hotel of Xinjiang University, participants will be housed in the Hotel of Huaqiao. However, one needs to take a walk about fifteen minites to get to the conference site.

The cost of one room Single 300 yuan/day (1 US Dollar=8.1 yuan about) Double 150 yuan/day each person 2). The meals duing the conference will be coverd by the conference.

Registration fees

The registration fee is 400 yuan, the registration fee will be used to refreshment, banquets, etc.

Registration You can register by 1). Complete the registration form and return it (together with the abstract

of your paper) to GUO Xiao-Feng by e-mail: xfguo@xju.edu.cn or fax: 86-991-2879377 or postal mail to: Prof. Xiaofeng Guo, Institute of Mathematics and Physics, Xinjiang Univer-

sity, Wulumuqi Xinjiang 830046, P. R. China.

2). or by e-mail to WEN Zhi-Ying: wenzy@tsinghua.edu.cn

Organizer Tsinghua University Xinjiang University

Sponsors NNSF of China Xinjiang University

Transports

There are flights from Beijing, Shanghai and Guangzhou to Urumqi every day, when you arrive at the airport of Urumqi, we will pick you up to the hotel. (When you have reserved your flight to Urumqi, please tell us the number of the flight by e-mail or fax).

Tourism

We will organize a tour in some place near Urumqi (for example, Tianchi Lake) during the conference, of which the expense will be covered by the conference. If there are enough people who want to take a tour in other place of Xinjiang after the conference, we will help you to contact some agent.

For example, a possible tour line is from Urumqi - Turpan(Tulufan) - Querle (Buositen lake) - Urumqi. It will spend 3 days (maybe August 27-29). The cost is about 1500 yuan each person.

REGISRATION FORM

1. PERSONAL INFORMATION

NAME:

(Family) (First)

SEX: NATIONALITY: MAILING ADDRESS:

CITY: STATE: COUNTRY: No. PASSPORT: FAX: TEL: E-MAIL:

2. HOTEL RESERVATION SINGLE: (about 300yuan/day) DOUBLE (about 150yuan/day each person) CHECK-IN DATE: CHECK OUT DATE: TOTAL NUMBER OF NIGHTS:

3. TOURISM Do you attend the tourism after the conference? Yes/no:

4. Please return this form as well as the abstract of your paper by e-mail (or by ordinary mail) to the one of:

Prof. GUO Xiao-Feng: xfguo@xju.edu.cn Prof. WEN Zhi-Ying: wenzy@tsing.hua.cn

Introduction of Xinjiang, Urumqi and South-Xinjiang

A survey of Urumqi

Being located at the northern foot of the Tianshan Mountains and on the edge of the Zhungar Basin, Urumqi, with an area of 49.3 sq-km, within the urban district and a population of about 1.2 million, is the capital of Xinjiang, in the past travellers on the Silk Road all had to pass through it.

The golden season of this city is from May to September, when flowers are all in blossom and various kinds of melons and fruits are everywhere in the city. In the suburbs of Urumqi there are many tourist spots and summer resorts such as the well-known Tianchi Lake, the densely wooded Nanshan Mountain, the stone coffins in the graves of the period of the Warring States in Wulabo, the ruins of an ancient city built in the Yuan Dynasty, etc. The traditional sports such as girl chasing, lamb tussing, horse race, etc. which are often held in the Nanshan pastureland will all the more attract the tourists.

A survey of Turpan

In the low-lying Turpan basin at the southern of the Tianshan Mountain in the central part at Xinjiang lies city of Turpan. The Turpan besin is the lowest area in China, in which the Aiding Lake, with the surface being 154m, below sea level, is the second lowest spot, next only to the Dead Sea in Jordan, it is also the hottest place in China, so called also the Fire Continent.

Turpan was called Gushi in ancient times. In the Han Dynasty it was first called the State Cheshiqian and later the State of Cheshiqianwang. In the Northern Wei Dynasty the Rourans established the Kindom of Gaochang in Turpan. In the Tang Dynasty Xichang Zhou was established here and later it was changed to Xi Zhou. In 1913, Turpan County was set up.

Turpan was an important and big town on the Silk Road. There are now in the suburbs of the Turpan City the ruins of some ancient cities built more than 1500 years ago and precious ”underground museums”-tombs of the ancient times from the 3rd to the 9th centuries.

A survey of Kashi

Kashi Prefecture is situated in the southern part of Xinjiang, bordering upon the Taklamaken Desert in the east and upon Afghanistan, Pakistan and India to the west and southwest. The Kashi city, is the capital of Kashi Prefecture, it is 1473 km. from Urumqi.

Kashi is the abbreviation for Kashgar. In 60 B. C., it was administered as Office of the Governor of the Western Regions. In the Han Dynasty, Ban Chao, with Kashi as his headquarters, stayed to built up the Western Regions for 30 years. After the establishment of Xinjiang in 1884, Shufu County was set up here. In 1952, part of the county was seperated and the Kashi City was established.

Kashi has an arid continental climate of the temperate zone with distinct seasons, of which the summer is long while the winter is short. The annual average is 11.7C and the average temperature of July is 25.9C.

Being an important open city on the ancient Silk Road, Kashi has always been a place where the eastern and western cultures are blended. In Kashi tourists can enjoy the Kashi Mukam. Kashi is a city with many scenic and historical sites as the Apahoja Tomb, the Etgar Mosque, the Yusup Hashajip Tomb, the ancient City of Isiksar, the ancient City of Ban Chao, the ruins of Genggongtai,etc.

Foreigners can tour Kashi at any time of the year, but the best season for tourism are summer and autumn.